A certain factory employed 600 men and 400 women and the average wage was rupees 25.50 per day. If a women got rupees 5 less than a man then what are their daily wages?

Grade:12th pass

3 Answers

Sehaj2122
6 years ago
20.50. 25.50- 5.00=20.50Are you a kid of 9 th grade or 1st.I think so.If you didn`t know answer of thisHow are u going to solve the difficult one.
Manisha Loomba
6 years ago
I am so sorry sir but your answer is wrong. . .thanks for your effort !If you have a problem in answering then please don`t answer but don`t make unnecessary comments
Nitesh
6 years ago
The daily wages of man and woman are rupee 28 and 23 respectively.Let wage of man is x.Then the wage of woman is x-5.Now the average wage is x+x-5 / 2 =25.50.and the x=28,x-5=28-5=23.So the wage of man is 28 and woman is 23.